# A LoD of Gaussians: Unified Training and Rendering for Ultra-Large Scale Reconstruction with External Memory

### v3

@@ 5.1 Results
OctreeGS 19.20 0.601 0.498 OOM OOM OOM
CityGS 21.40 0.692 0.383 19.47 0.645 0.531
H-3DGS (leaves) 24.35 0.788 0.273 17.84 0.6101 0.4667
Ours 24.61 0.774 0.297 21.85 0.707 0.456
(a) Ground Truth
(b) Ours
(c) CityGaussian
(d) Hierarch. 3DGS
(e) HorizonGS
Figure 6: Qualitative comparison of our method and SOTA methods (CityGaussian; Hierarchical; OctreeGS) on the MC-small-city+ and campus datasets.
@@
Our level-of-detail and caching strategy can also be applied to efficiently render the finished models: As supplemental material, we include fly-through videos of all scenes rendered on an RTX 3090 GPU at this anonymized link: https://www.youtube.com/playlist?list=PLocO6QqSKZ4YN7DyuC5gWEeVciSV6RdAk. Table 1 demonstrates effective VRAM reduction during rendering of test images in comparison with other large-scale 3DGS methods. Comparisons of rendering VRAM with 3DGS (3DGS) and gsplat (Gsplat) are included in Appendix A.4.1.
6 Discussion and Outlook
A LoD of Gaussians enables seamless training and rendering of ultra-large 3DGS models on consumer hardware. By storing Gaussian data in external memory and streaming it on demand, our method avoids the pitfalls of chunk-based pipelines. The hierarchical SPT structure accelerates LoD selection and remains robust to ongoing training changes. Combined with caching and view selection, our system significantly reduces out-of-core overhead. Together, these components enable efficient reconstruction and rendering at scale, demonstrated on the MC-small-city+ dataset.
Limitations and Future Work
A primary limitation of our method lies in initialization. Accurate estimation of camera poses and sparse point clouds remains challenging at large scale, particularly for real-world datasets with sparse or inconsistent coverage. Future work could explore improved initialization strategies or joint optimization of poses and Gaussians during early training.
Our system also requires roughly 1 GB of RAM per million Gaussians, which—while more efficient than prior methods—still constrains scalability. Loading from disk is feasible in our experiments, though at the cost of a $10\times$ slowdown. Further, while our method in general greatly reduces the number of training iterations necessary, individual training iterations can take a lot longer than standard 3DGS due to the loading overhead.
The level-of-detail system makes our method robust against vast differences in view distance, but when these differences are not present, like in single-height aerial datasets as evaluated in Appendix A.4, it becomes a burden instead, making more straight-forward divide-and-conquer training the better choice.
Interactive rendering could be further optimized by avoiding per-frame hierarchy cut recomputation and enabling asynchronous streaming. While frustum culling reduces memory load in most views, it is ineffective when the entire scene falls inside the frustum. Occlusion culling could address this by skipping entire SPTs before loading.
Overall, we believe that out-of-core 3D Gaussian Splatting is a promising direction for scaling radiance field methods to city-scale and beyond, without the need for specialized hardware.
Reproducibility Statement
To ensure reproducibility of our results, we have appended our entire codebase along with the config files of our method and baselines required to train each of the scenes. Detailed breakdowns of datasets and hyperparameters can be found in Appendix A.4 and A.5.
Appendix A Appendix
A.1 Figures
(a) Divide and Conquer
(b) Ours
Figure 7: Comparison between the divide and conquer training process as used in Hierarchical; VastGaussian; CityGaussian; HorizonGS to our training process. Colored regions are present in VRAM, training views are drawn in red.
[…]
Figure 9: Frustum Culling and LoD selection (left) greatly reduces the number of Gaussians required to render a view (right).
Figure 10: Hierarchical SPTs enable smooth transitions between detailed (left) and coarse (right) representation.
(a) Chunk Bleeding in H-3DGS (Hierarchical).
(b) Chunk Ghosting in CityGaussian (CityGaussian).
Figure 11: Artifacts caused by the divide-and-conquer strategy on the MC-small-city+ dataset.
Figure 12: Training image rendered during the iteration depicted in Figure 5.
Figure 13: Images from H-3DGS on the MatrixCity-Scale scene are obscured by large floaters.
Figure 14: All Gaussian properties except gradients are always present in CPU RAM, whereas only the currently loaded Gaussians and slim SPT information needs to be stored on GPU.
Figure 15: SPTs for a frame of MatrixCity rendered in different colors.
Figure 16: A LoD of Gaussians is able to reconstruct and render huge environments with 150M+ Gaussians with large variations of scale on consumer hardware.
(a) Image generated during chunk training by H-3DGS.
(b) The same image generated during training by A LoD of Gaussians.
Figure 17: Chunk-based optimizations causes floaters to expand outside of the chunk boundary.
A.2 Initialization and Training
Following Hierarchical, we initialize the Gaussian model from a sparse point cloud, augmented with skybox points. This initial representation is small enough to fully reside in GPU memory and is trained for 100100k iterations without densification. The goal of this phase is to establish a stable global scene structure before constructing the hierarchy. After this initial optimization, we build a binary Gaussian hierarchy, where the trained Gaussians act as leaf nodes and the parent nodes represent merged approximations of their two children.
The Gaussian properties–including base color, SH coefficients, position, and covariance–are optimized using the standard loss propagation introduced by 3DGS. To fulfill our tight memory requirements, we update only the parameters of the cut set chosen for each training view. Thus, gradients may, in general, be propagated to Gaussians in the middle of the hierarchy. Assuming good coverage and diversity of training views, these updates will diffuse over the entire hierarchy over the course of training, leading to smooth transitions from the highest to the lowest LoD (see Figure 10).
A.3 Ablations
@@
We assess the contribution of key components through an ablation study, using recorded camera paths across all scenes (see supplemental videos). Table 3 reports average frame times over these paths. Caching significantly improves performance, roughly doubling the framerate across scenes. Frustum culling is particularly beneficial for large-scale models but may incur overhead in smaller scenes. To evaluate cut efficiency, we compare the time required to compute the visible set using either full hierarchy BFS or our HSPT-based approach. HSPT consistently yields faster cut times. For training, we measure average iteration durations over 1 000 steps. Here, frustum culling and caching prove essential, substantially reducing the number of Gaussians loaded and processed per view.
@@
Table 3: Ablations. Average frame times for rendering camera paths and average iteration times during training with and without caching Gaussians. The final results show the average timings of the hierarchy cut during rendering using our HSPT and the baseline BFS approach. For campus, we evaluate two differently sized models with 38M and 80M Gaussians respectively.
@@
Render w/o Frustum Culling 52.3 ms 89.3 38.3 ms 110.2 ms
Train Full 156 ms 174 ms 205 ms
Train w/o Cache 471 ms 257 ms 244 ms
Train w/o Frustum Culling 685 ms 626 ms 312 ms
HSPT Cut (Render) 31.9 ms 45.7ms 31.3 ms 36.5 ms
BFS Cut (Render) 47.8 ms OOM 40.0 ms 53.7 ms
@@
When training MC-smaller-city+ dataset (250k iterations, up to 60M Gaussians), enabling our guided view selection reduces the number of Gaussians loaded from RAM per frame from 355 770 to 219 043—a $35\%$ reduction in memory transfers. Reconstruction quality is unaffected, with PSNR and SSIM improving slightly by 0.04 and 0.05, respectively.
@@
To evaluate the effectiveness of our level-of-detail system, we choose 5 random images from each dataset and compare the resulting quality metrics and number of Gaussians rendered at 50 different levels of details. The results are visualized in Figure 18. In the vast majority of cases, increasing the LOD level consistently improves rendering quality. Overall, we only notice a very small dip in quality when reducing the Gaussian count by 50%, verifying the effectiveness of our continuous LoD structure.
(a) MC-small-city+ aerial
(b) MC-small-city+ street
(c) campus
Figure 18: Quantitative evaluation of our LoD system for 5 randomly chosen views and 50 different levels of detail.
A.4 Dataset Details and Further Results
@@
For our main benchmark, MC-small-city+, we aggregate 33 006 street-view images and 7 672 aerial views from the small-city scene of the MatrixCity dataset (MatrixCityDataset) and add 533 additional high-altitude views we generated manually. We evaluate reconstruction quality on a separate set of 4 228 test views, according to the test split provided by the dataset. The scene is extremely challenging due to its enormous scale, sparse views and wide variation in scale: As such, we use a subset of the entire dataset (15.1k images, covering about a third of the area), denoted as MC-smaller-city+ for baselines that are not able to reconstruct the full dataset. For MC-small-city+, we use the camera poses provided by the MatrixCity dataset and convert them to the COLMAP format for Gaussian splatting. We merge the provided street and aerial sparse point clouds and randomly downsample them by a factor of $5$ to get more realistic initialization conditions. We do not make use of the ground truth depth images provided by the dataset, as we consider this to be an unrealistic advantage. Because the dataset does not provide point correspondences required for scaling monocular depth maps, we deactivate depth supervision in our method and baselines for this dataset. Figure 19 shows the scene along with the camera distribution. While some methods have successfully reconstructed only aerial (CityGaussian; CityGaussianV2; OctreeGS) or only street-level views (CityGaussianV2; NerfXL), training a model that holds up to scrutiny from both perspectives presents a particular challenge. Training on close and far views simultaneously–without a proper LoD system like ours in place–significantly degrades visual quality for both, as noted by HorizonGS; CrossView. Moreover, such a scenario complicates partitioning the scene into independent chunks.
@@
Figure 19: The MC-small-city+ dataset spans multiple city blocks, which are supervised by tens of thousands of views, varying widely in scale.
@@
H-3DGS Hierarchical did not use a test split for their dataset, but we hold back every $100$th image alphabetically for testing. We reevaluate H-3DGS for this scene with the new test split, using the camera calibrations and chunk splits provided on their website. In accordance with the instructions, we disable exposure optimization for evaluation. On this dataset, we use depth supervision identical to Hierarchical.
@@
To demonstrate our ability to reconstruct small scenes, we also evaluate our method on the smaller, single-chunk versions of the Campus and Small City (cf. Table 4). The quality metrics indicate that training on the LoD structure (which would not be required for scenes of this scales) during training only leads to marginal reductions in visual quality. The distribution of street views in the campus dataset is visualized in Figure 20(b).
Table 4: H-3DGS single chunk view synthesis results. Results reproduced from Hierarchical.
@@
Mill19 dataset
Table 5 shows results on the Mill19 dataset (MegaNerf). We use the camera poses and sparse point cloud provided by CityGaussian. In accordance with baselines, we downscale all images by a factor of 4. We have chosen this dataset, because it is widely used in large-scale novel-view synthesis. At the same time, this dataset represents a worst-case scenario for our method, as the regular, same-height aerial views (cf. Figure 20(a)) negate any benefit of our level-of-detail method and can be split into independent chunks trivially.
Table 5: Mill19 novel view synthesis results Results with are reproduced from CityGaussian. Results of H-3DGS are reproduced from Hierarchical.
@@
Figure 20: Overview of the aerial and street datasets.
A.4.1 Rendering Comparison
To demonstrate the effectiveness of our LoD method, we compare VRAM usage and image quality to rendering the highest level of detail leaf Gaussians with both 3DGS (3DGS) and gsplat (Gsplat) using their default rasterization settings. We conducted comparisons across three different viewpoints to assess performance at varying levels of detail. As illustrated in Figure 21, our approach achieves visual fidelity comparable to the baselines while significantly reducing VRAM usage, highlighting the effectiveness of our LOD scheme.
(a) Overview
(b) Aerial View
(c) Street View
Figure 21: Comparision to standard 3DGS (3DGS) and GSplat (Gsplat) in terms of rendering quality and VRAM usage.
A.5 Evaluation Details
Finding baselines for the MC-small-city+ scene proved challenging, as the only methods we are aware of that have successfully trained on all fused aerial and street-level views–RetinaGS and NerfXL–both require 64 GPUs running in parallel. We were not able to meet these hardware requirements for evaluation and hold that these methods and similar multi-GPU works like Grendel are orthogonal to our method. H-3DGS (Hierarchical) has demonstrated the ability to reconstruct similar sized datasets, but its evaluation was restricted to street-level views. We choose Hierarchical and other popular large-scale Gaussian Splatting methods (CityGaussian; HorizonGS; OctreeGS) as baselines. Note that ”large-scale” is often used ambiguously in the context of novel-view synthesis, alternatively meaning reconstructing very complex scenes with large number of images or reconstructing novel views using far-away (i.e. aerial) input views: With the exception of Hierarchical, the biggest datasets these methods use in their papers consist of just 5620 images. In order to train MC-small-city+ with 42.2 thousand images, it was necessary to modify CityGaussian, HorizonGS and OctreeGS to load images from disk instead of caching them in RAM or VRAM. This slows down training, but does not affect the final results.
Most large-scale 3DGS frameworks have complicated multi-stage training processes and are very sensitive to hyperparameters. We consider it a benefit of our training pipeline that the user is not confronted with tuning the error-prone partitioning process, being more akin to the original Gaussian Splatting training. In general, the hyperparameters suggested for large-scale scenes were used for baselines. If provided, we use configurations explicitly targeting the MatrixCity dataset, if it improves metrics over default configuration. Learning rates were not modified, as they were similar to ours and should not be a barrier to reconstruction on this scale. We also reached out to the authors of all our baselines about suggesting hyperparameters for the MC-small-city+ dataset, which we incorporated when provided. For methods that make use of configuration files, we include them in the supplemental material.

### v4

@@ 5.1. Results
Dataset Hierarchical 3DGS OccluGaussian
Scene Small City (5.8k) Campus (22.0k) Canteen (8.6k) Classbuilding (9.1k)
[…]
OctreeGS 19.20 0.601 0.498 20.54 0.643 0.544 21.49 0.826 0.293 23.08 0.889 0.197
CityGS 21.40 0.692 0.383 19.47 0.645 0.531 20.41 0.794 0.275 20.48 0.840 0.244
H-3DGS 24.35 0.788 0.273 17.84 0.610 0.467 22.71 0.825 0.178 23.87 0.881 0.128
CLM-GS 22.54 0.699 0.392 16.60 0.613 0.558 23.13 0.864 0.168 22.89 0.892 0.159
OccluGS N/A N/A N/A N/A N/A N/A 25.25 0.900 0.100 25.33 0.921 0.083
Ours 24.61 0.774 0.297 21.85 0.707 0.456 23.33 0.867 0.155 24.14 0.915 0.116
@@
Our level-of-detail and caching strategy can also be applied to efficiently render the trained models. As supplemental material, we include fly-through videos of the evaluated scenes rendered on an RTX 3090. Table 1 demonstrates effective VRAM reduction during rendering of test images compared to other 3DGS methods. Figure A LoD of Gaussians: Out-of-Core Training and Rendering for Seamless Ultra-Large Scene Reconstruction compares rendering VRAM consumption and image quality of our LoD method with full-detail 3DGS (Kerbl et al., 2023) and gsplat (Ye et al., 2025). Our approach achieves visual fidelity comparable to the baselines while significantly reducing VRAM usage, highlighting the effectiveness of our LoD scheme.
Ablations
@@
We assess the contribution of key components through an ablation study using recorded camera paths across a selection of scenes (see supplemental videos). Table 3 reports average frame times over these paths. Caching significantly improves rendering performance, roughly doubling the framerate across scenes by reducing the average number of Gaussians loaded from RAM by 93% on Campus and 86% on MC-small-city+. The effectiveness of frustum culling scales with scene size: On the full MC-small-city+ scene, 24.5 million Gaussians are frustum culled on average (88% reduction), while for Campus and Small City the corresponding values are 9.8 million and and 7.9 million (74% and 65% reduction), respectively. As expected, the overhead of frustum culling amortizes with scene size, making it essential for the largest datasets.
To evaluate cut efficiency, we compare the time required to compute the visible set using either full hierarchy BFS or our HSPT-based approach. HSPT consistently yields faster cut times due to improved parallelization. Moreover, the BFS approach requires positions and scales for all Gaussians to reside in memory, causing it to exceed 24GB of VRAM on MC-small-city+, whereas the HSPT method peaks at 21GB. For training ablations, we measure average iteration durations over 1 000 steps. Here, frustum culling and caching prove essential, substantially reducing the number of Gaussians loaded and rendered per view.
@@
Table 3. Ablations. Average frame times for rendering camera paths and average iteration times during training with and without caching Gaussians. The final results show the average timings of the hierarchy cut during rendering using our HSPT and the baseline BFS approach. For Campus, we evaluate two different models with 38M and 80M Gaussians respectively.
@@
Render w/o Frustum Culling 52.3 ms 89.3 ms 38.3 ms 110.2 ms
Training Full 156 ms 174 ms 205 ms
Training w/o Cache 471 ms 257 ms 244 ms
Training w/o Frustum Culling 685 ms 626 ms 312 ms
Render Cut (HSPT) 31.9 ms 45.7ms 31.3 ms 36.5 ms
Render Cut (BFS) 47.8 ms OOM 40.0 ms 53.7 ms
6. Discussion and Outlook
A LoD of Gaussians enables seamless training and rendering of ultra-large 3DGS models on consumer hardware. By storing Gaussian data in external memory and streaming it on demand, our method avoids the pitfalls of chunk-based pipelines. The HSPT datastructure accelerates LoD selection and remains robust to ongoing training changes. Combined with caching and view selection, our approach significantly reduces out-of-core overhead. These components enable efficient reconstruction and rendering at scale, as demonstrated on challenging multi-scale scenes such as MC-small-city+.
Limitations and Future Work
Our method represents an informed trade-off between performance and memory. While it greatly reduces the necessary number of training iterations for large-scale scenes, individual iterations take longer than standard 3DGS training due to data loading and hierarchy management overhead.
Similarly, although rendering framerates are a significant improvement over neural-based and other out-of-core methods, they do not yet reach the peak performance of fully in-core 3DGS.
Our approach requires roughly 1 GB of RAM per million Gaussians, which—while more efficient than prior methods—still constrains scalability for extremely large scenes. Loading from disk is feasible in our experiments, but at the cost of about a $10\times$ slowdown, making fast secondary storage highly desirable.
The level-of-detail system makes our method robust to large variations in view distance. However, when such variation is absent—for example in single-height aerial datasets as evaluated in Suppl. A.3—the LoD machinery introduces unnecessary overhead, making more straightforward training a competitive alternative. Interactive rendering performance could be further improved by avoiding per-frame hierarchy cut recomputation and asynchronous streaming.
While frustum culling effectively reduces memory load in most views, it is ineffective when the entire scene lies inside the frustum. Occlusion culling could address this limitation by skipping entire SPTs before they are loaded into memory.
Overall, we believe that out-of-core 3D Gaussian Splatting is a promising direction for scaling radiance field methods to city-scale scenes and beyond on consumer-grade hardware.
Appendix A Supplementary Material
Our full supplementary material consists of:
  1. (1)
    This document, which includes additional ablation studies, implementation details and experiments, which we could not include in the main paper due to page restrictions
  2. (2)
    A short video presentation about this manuscript
  3. (3)
    An html-page with more qualitative comparisons and short videos
  4. (4)
    Our full source code contained in the Code directory
  5. (5)
    Training configurations detailing all hyperparameters for our and baseline methods on each experiment
A.1. Additional Ablations
We present additional experiments to demonstrate the effectiveness of each component of our method.
@@ A.1. Additional Ablations
To verify the effectiveness of caching and view selection, we conduct an experiment by training the MC-smaller-city+ dataset (100k iterations, up to 30M Gaussians) with and without view selection and caching. The results in Table 4 show that while quality metrics are barely affected (with a maximal difference of 0.006 PSNR), the number of Gaussians that need to be loaded into VRAM decrease by a factor of almost $20\times$ with view selection and caching enabled.
Table 4. Ablation results on the MC-smaller-city+ dataset. #Loaded refers to the average number of Gaussians loaded from RAM over 100 000 iterations.
[…]
w/o caching 21.831 0.701 0.425 1 629 187
w/ caching 21.829 0.700 0.429 290 572
w/ view selection 21.825 0.697 0.432 89 451
@@
To evaluate the effectiveness of our level-of-detail system, we choose 5 random images from each dataset and compare the resulting quality metrics to the number of Gaussians rendered at 50 different levels of details. The results are visualized in Figure 16 in addition to average FPS for rendering the view 10 times with pre-warmed cache on an H200 GPU. As expected, FPS scales closely with the inverse of the number of Gaussians. In the vast majority of cases, increasing the LoD level consistently improves rendering quality. Overall, we only notice a small dip in quality when reducing the Gaussian count by 50%, verifying the effectiveness of our LoD structure. Further, using continuous instead of discrete levels of detail leads to the smoothness of the curves in Figure 16, indicating a smooth transition between LoD levels, which reduces popping.
Effectiveness of Caching
In Figure 15, we provide sensitivity curves for cache size vs. FPS during rendering. In each iteration up to cache size Gaussians from the previous frames may be reused. The results once again demonstrate the importance of the caching system to rendering efficiency. The graphs show that performance increases steeply with increasing cache size until it exceeds the typical number of Gaussians required per iteration and starts to plateau.
(a) MC-small-city+
(b) Campus
Figure 15. Average FPS for rendering 1000 frames of the camera paths for the MC-small-city+ and Campus scenes with various cache sizes.
(a) MC-small-city+ aerial.
(b) MC-small-city+ street.
(c) Campus.
Figure 16. Qualitative evaluation of our LoD system for 5 randomly chosen views and 50 different levels of detail.
A.2. Initialization and Training Details
Following Kerbl et al. (2024), we initialize the Gaussian model from a sparse point cloud, augmented with skybox points. This initial representation is small enough to fully reside in GPU memory and is trained for 100100k iterations without densification. The goal of this phase is to establish a stable global scene structure before constructing the hierarchy. After this initial optimization, we build a binary Gaussian hierarchy, where the trained Gaussians act as leaf nodes and the parent nodes represent merged approximations of their two children.
The Gaussian properties–including base color, SH coefficients, position, and covariance–are optimized using the standard loss propagation introduced by Kerbl et al. (2023). To fulfill our tight memory requirements, we update only the parameters of the cut set chosen for each training view. Thus, gradients may, in general, be propagated to Gaussians in the middle of the hierarchy. Assuming good coverage and diversity of training views, these updates will diffuse over the entire hierarchy over the coarse of training, leading to smooth transitions from the highest to the lowest LoD.
A.3. Dataset Details and Further Results
@@
For our main benchmark, MC-small-city+, we aggregate 33 006 street-view images and 7 672 aerial views from the small-city scene of the MatrixCity dataset (Li et al., 2023) and generate 533 additional high-altitude views. We evaluate reconstruction quality on a separate set of 4 228 test views, according to the test split provided by the dataset. The scene is extremely challenging due to its enormous scale, sparse views and wide variation in scale: As such, we also construct a subset of the entire dataset (15.1k images, covering about a third of the area), denoted as MC-smaller-city+ for baselines that are not able to reconstruct the full dataset. For MC-small-city+, we use the camera poses provided by the MatrixCity dataset and convert them to the COLMAP format for Gaussian splatting. We merge the provided street and aerial sparse point clouds and randomly downsample them by a factor of $5$ to get more realistic initialization conditions. We do not make use of the ground truth depth images provided by the dataset, as we consider this to be an unrealistic advantage. Figure 17 shows the scene along with the camera distribution. While some methods have successfully reconstructed only aerial (Liu et al., 2024a, b; Ren et al., 2024) or only street-level views (Liu et al., 2024b; Li et al., 2024b), training a model that holds up to scrutiny from both perspectives presents a particular challenge. Training on close and far views simultaneously–without a proper LoD system like ours in place–significantly degrades visual quality for both, as noted by Jiang et al. (2025); Zhang et al. (2024). Moreover, such a scenario greatly complicates partitioning the scene into independent chunks.
Why are the quality metrics on MC-small-city+ so low?
When comparing our evaluation on the MatrixCity dataset with those in other works like Liu et al. (2024a); Zhao et al. (2026); Jiang et al. (2025); Ren et al. (2024), it might seem surprising that they achieve significantly lower quality metrics on our constructed MC-small-city+ scene. There are several reasons for this:
  1. (1)
    These works only evaluate on the aerial views, which make up less than 20% of the scene images and are generally easier to reconstruct.
  2. (2)
    Reconstructing street, aerial and the newly generated high-aerial views in the same scene is much more difficult than reconstructing just aerial views.
  3. (3)
    Some competing works (Liu et al., 2024a; Zhao et al., 2026) use the ground-truth depth maps or dense point clouds from the MatrixCity dataset. We would consider this to be an unrealistic advantage, since they contain almost the entire ground truth geometry of the scene. Zhao et al. (2026) even starts the reconstruction from the full dense point cloud and never densifies, whereas our evaluation starts from the same point cloud downsampled by a factor of 5.
@@
Figure 17. The MC-small-city+ dataset spans hundreds of buildings, which are supervised by tens of thousands of views, varying widely in scale.
Uni10K
In order to address the gap of real-world large-scale datasets that are captured across multiple scales, we present Uni10k. The scene consists of an outdoor campus of approximately $100\,000\text{ m}^{2}$, captured from both ground-level and aerial perspectives. Standard reconstruction via COLMAP at this scale would typically require weeks of computation due to the complexity of the image-matching and mapping stages. To mitigate this, we leverage spatial and temporal priors in conjunction with a coarse-to-fine scheme. Specifically, we utilize GPS data to reduce the matching complexity from quadratic to near-linear by limiting image comparisons to a predefined spatial radius. Regarding temporal priors, since frames are sampled from video sequences, we initially reconstruct a baseline model starting from a set of frames uniformly sampled every second of video. Then, we densify the camera coverage by incrementally registering, triangulating, and refining new images using only local BA. The process concludes with several rounds of Global bundle adjustment. The final reconstruction comprises more than 10 00010\,000 images at 4K resolution and  6.2M sparse points, with an overall MRE of approximately $0.63$ pixels. We train on the full resolution images and hold back every 8th frame alphabetically for the test set. Figure 18 shows the sparse point cloud and camera distribution.
Figure 18. Uni10K’s sparse point cloud. The scene is reconstructed from more than $10$ thousand 4k images at widely varying heights, both from aerial and ground perspective.
@@
Since H-3DGS Kerbl et al. (2024) did not introduce a test split for their dataset, we hold back every $100$th image alphabetically for testing. We reevaluate H-3DGS for this scene with the new test split, using the camera calibrations and chunk splits provided on their website. In accordance with the instructions, we disable exposure optimization for evaluation. On this dataset only, we use depth supervision identical to Kerbl et al. (2024). The distribution of street views in the Campus scene is visualized in Figure 19(b).
@@
To demonstrate our ability to reconstruct small scenes, we also evaluate our method on the smaller, single-chunk versions of the Campus and Small City scenes (cf. Table 5). The quality metrics indicate that training on the LoD structure (which would not be required for scenes of this scales) during training only leads to marginal reductions in visual quality.
Table 5. H-3DGS single chunk view synthesis results. Results with are taken from (Kerbl et al., 2024).
@@
OccluGaussian Dataset
The OccluGaussian dataset features three large-scale indoor environments. Unfortunately, the download link for the Gallery scene on the official website leads to a broken zip file. As we have not received an answer from the authors, we were unable to evaluate on this scene. Additionally, some training images that are referenced in the camera pose files were not included in the download of the Canteen scene. We ran the evaluation on all images that were provided using full resolution images.
Mill19 and Urbanscene3D dataset
Table 6 and 7 show results on the Mill19 (Turki et al., 2022) and UrbanScene3D (Lin et al., 2022) datasets. We use the camera poses and sparse point cloud provided by Liu et al. (2024a). In accordance with baselines, we downscale all images by a factor of 4. We have included these dataset, because they are widely used in large-scale novel-view synthesis. At the same time, they represents a worst-case scenario for our method, as the regular, same-height aerial views (cf. Figure 19(a)) negate any benefit of our level-of-detail method and can be split into independent chunks trivially.
Table 6. Mill19 novel view synthesis results Results with are taken from Liu et al. (2024a). Results of H-3DGS are taken from Kerbl et al. (2024).
@@
Table 7. UrbanScene3D novel view synthesis Results. Results with are taken from Liu et al. (2024a).
Residence (2.5k images) Sci-Art (3k images)
[…]
CityGaussian 21.90 0.805 0.217 21.34 0.833 0.232
3DGS 21.44 0.791 0.236 21.05 0.830 0.242
H-3DGS 17.52 0.416 0.391 18.42 0.652 0.340
OctreeGS 18.76 0.523 0.539 18.97 0.620 0.535
Ours 20.25 0.705 0.319 20.70 0.794 0.272
@@
Figure 19. Overview of the aerial and street datasets.
A.4. Evaluation Details
Finding baselines for the MC-small-city+ scene proved challenging, as the only methods we are aware of that have successfully trained on all fused aerial and street-level views–Li et al. (2024a) and Li et al. (2024b)–both require 64 GPUs running in parallel. We were not able to meet these hardware requirements for evaluation and hold that these methods and similar multi-GPU works like Zhao et al. (2024) are orthogonal to our method. H-3DGS (Kerbl et al., 2024) has demonstrated the ability to reconstruct similar sized datasets, but its evaluation was restricted to street-level views. We choose Kerbl et al. (2024) and other popular large-scale Gaussian Splatting methods (Liu et al., 2024a; Jiang et al., 2025; Ren et al., 2024; Zhao et al., 2026) as baselines. Table 8 compares the supported features of our method and all baselines.
Most of the large-scale 3DGS frameworks have complicated multi-stage training processes and are very sensitive to hyperparameters, which we detail in the following section. We consider it a benefit of our training pipeline that the user is not confronted with tuning the error-prone partitioning process, being more similar to the original Gaussian Splatting training.
LoD Training LoD Rendering Out-of-Core Training Interactive Rendering No Chunking Artifacts Gaussians
CityGaussian Explicit
HorizonGS Neural
H-3DGS Explicit
OctreeGS Neural
CLM-GS Explicit
VastGaussian Explicit
OccluGaussian Explicit
Ours Explicit
Table 8. Comparison of features.